OpenStudy (anonymous):

How many solutions are there to the following system of equations? 6x + 12y = 11 4x + 8y = 3 (Points : 4) 0 1 2 infinitely many

OpenStudy (solomonzelman):

6x + 12y = 11 --> divide by 6 ---> x+2y=11/6 4x + 8y = 3 --> divide by 4 ---> x+2y=3 Does 3 = 11/6 ?

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

its 0 no real solution

OpenStudy (solomonzelman):

YES!

OpenStudy (anonymous):

Solve the following system of equations by linear combination: 2d – e = 7 d + e = 5 (Points : 4) There is no solution. There are an infinite number of solutions. The solution is (3, –1). The solution is (4, 1).

OpenStudy (solomonzelman):

No solution. Reasoning. x+2y=11/6 also we know that x+2y=3 so if x+2y, =11/6 and =3, THEN 3=11/6 which is not true.

OpenStudy (solomonzelman):

2d – e = 7 d + e = 5 subtract them from each other d=2 2+e=5, e=3 d=2 e=3